1/2

Christian Morgenstern, Christian Morgenstern,  Christian Morgenstern,

Der Werwolf The Banshee (An Approach)

The Werewolf

tradukita de Max Knight  tradukita de Alexander Gross

Ein Werwolf eines Nachts

entwich von Weib und Kind und sich

begab an eines Dorfschullehrers

Grab und bat ihn: �Bitte, beuge

mich!“ One night, a banshee slunk

away from mate and child, and in

the gloom went to a village teacher’s

tomb, requesting him: “Inflect me,

pray.”

A Werewolf, troubled by his name, Left wife and brood one night and

came To a hidden graveyard to enlist The aid of a long-dead philologist.

Der Dorfschulmeister stieg

hinauf auf seines Blechschilds Mes-

singknauf und sprach zum Wolf, der

seine Pfoten geduldig kreuzte vor dem

Toten:

The village teacher climbed

up straight upon his grave stone with its

plate and to the apparition said who meekly knelt before the

dead:

“Oh sage, wake up, please don’t be-

rate me,” He howled sadly, “Just conjugate

me.” The seer arose a bit unsteady Yawned twice, wheezed once, and

then was ready.

�Der Werwolf“ – sprach der

gute Mann, �des Werwolfs, Genitiv so-

dann, dem Werwolf, Dativ, wie

man’s nennt, den Werwolf, - damit hat’s

ein End!“ “The banSHEE, in the sub-

ject’s place; the banHERS, the possessive

case. The banHER, next, is what

they call objective case–and that is all.”

“Well, ‘Werewolf’ is your plural

past, While ‘Waswolf’ is singularly cast: There’s ‘Amwolf’ too, the present

tense, And ‘Iswolf,’ ‘Arewolf’ in this same

sense.”

Dem Werwolf schmeichelten

die Fälle er rollte seine Augenbälle. �Indessen“, bat er, �füge

doch zur Einzahl auch die Mehr-

zahl noch!“ The banshee marveled at the

cases and writhed with pleasure,

making faces, but said: “You did not add, so

far, the plural to the singular!”

“I know that–I’m no mental

cripple– The future form and participle Are what I crave,” the beast rep-

lied. The scholar paused–again he tried:

…  …  …

2/2

Der Dorfschullehrer aber

musste gestehn, dass er von ihr

nichts wusste. Zwar Wölfe gab’s in großer

Schar, doch �Wer“ gab’s nur im Si-

nular.

The teacher, though, admit-

ted then that this was not within his

ken. “While bans are frequent”, he

advised, “A she cannot be plurized.”

“A ‘Will-be-wolf?’ It’s just too

long: ‘Shall-be-wolf?’ ‘Has-been-wolf?’

Utterly wrong! Such words are wounds beyond all

suture– I’m sorry, but you have no future.”

Der Wolf erhob sich

tränenblind – er hatte ja doch Weib und

Kind! Doch da er keine Gelehrter

eben, so schied er dankend und er-

geben.

The banshee, rising clammily, wailed: “What about my fami-

ly?” Then, being not a learned

creature, said humbly “Thanks” and left

the teacher.

The Werewolf knew better–his

sons still slept At home, and homewards now he

crept, Happy, humble, without apology For such folly of philology.

Verkinto de tiu ĉi Germana poemo estas Christian Morgenstern (Kristiano Matenstelo, ∗1871-05-06 – †1914-03-31).

Arg-825-1652 (2013-02-11 18:44:21)

Traduko de la Germana poemo “Der Werwolf” de Christian Morgenstern (Kristiano Matenstelo, ∗1871-05-06 – †1914-03-31) en la Anglan de Max Knight (civila nomo: Max Kühnel, ∗1909-06-08 – †1993-08-31).

Arg-825-1660 (2013-02-13 22:36:06)

Tiu ĉi angligo troviĝas en http: // bertilow. com/ literaturo/ lupfantomo. html . Pri la tradukinto Max Knight vidu: http: // www. jbeilharz. de/ morgenstern/ morgenstern_poems. html kaj http: // www. 123people. de/ ext/ frm? ti= ersonensuche% 20telefonbuch&search_ term= m x% 20k% C3% BChnel&search_country= DE&st= suc e% 20nach% 20personen&target_url= aHR0cDovL d3dy5saXRlcmF0dXJoYXVzLmF0L2luZGV4LnBocD9pZD02ODc4 .

Traduko de la Germana poemo “Der Werwolf” de Christian Morgenstern (Kristiano Matenstelo, ∗1871-05-06 – †1914-03-31) en la Anglan de Alexander Gross.

Arg-825-1657 (2013-02-13 18:39:45)

Tiu ĉi angligo troviĝas en http: // bertilow. com/ literaturo/ lupfantomo. html .

http://www.poezio.net/version?poem-id=825&version-id=,1652,1660,1657